Kamala borrowed ₹ 26400 from a Bank to buy a scooter at a rate of 15% p.a. compounded yearly. What amount will she pay at the end of 2 years and 4 months to clear the loan? (Hint: Find A for 2 years with interest is compounded yearly and then find SI on the 2nd year amount for 𝟒 /𝟏𝟐 years.) 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Principal (P) =₹26,400 Rate (R) = 15% per annum
Number of years (n) =2 whole 4/12
The amount for 2 years and 4 months can be calculated by first calculating the amount for 2 years using the compound interest formula, and then calculating the simple interest for 4 months on the amount obtained at the end of 2 years.
First, the amount for 2 years has to be calculated.
A =P (1 +R/100 )^n
=₹ [26400(1+ 15/100 )^2]
=₹ [26400 × (3/20)^2]
=₹ [26400 (23/20)^2]
=₹ 34914
By taking₹34,914 as principal, the S.I. for the next1/3years will be calculated.
S.I. = 34914 x 1/3 x 15
100
=₹1745.70
Interest for the first two years =₹ (34914 − 26400) = ₹8,514
And interest for the next1/3year =₹1,745.70
Total C.I. =₹(8514 +₹1745.70) =₹10,259.70
Amount = P + C.I. =₹26400 +₹10259.70 =₹36,659.70

​Thanks